OpenStudy (anonymous):

Find the Concavity: f(x) = x - 2cosx on the closed interval [0, 2pi]

OpenStudy (anonymous):

I know I need the second derivative f'(x) = 1+2sinx f''(x) = 2cosx Where do I go from there?

OpenStudy (campbell_st):

ok... so look at the stationary points let f'(x) = 0 and solve..... there are 2 stationary points 1 + 2sin(x) = 0 the solutions are exact values in 3rd and 4th quadrants... test them in the 2nd derivative... you'll also need to solve the 2nd derivative as the curve will have a point of inflection, given it has 2 stationary points... hope it helps

OpenStudy (anonymous):

Sorry I'm here

OpenStudy (anonymous):

I already found that Min = -2 and Max = 2pi-2 Is that what you mean?

OpenStudy (anonymous):

no you need to find f''(x) = 0

OpenStudy (anonymous):

Well cos = 0 at pi/2 and 3pi/2 within 0 to 2pi Is that it?

OpenStudy (campbell_st):

well not quite you are solving the 1st derivative \[0 = 1 + 2\sin(x)\] now solve for x..... this will give the the x values of any stationary points.... and there are 2 of those...

OpenStudy (campbell_st):

there are also 2 points of inflection in the domain, so there are 2 concavities

OpenStudy (anonymous):

-1 = 2sin(x) -1/2 = sinx But that is not within [0,2pi]

OpenStudy (anonymous):

cos(x) = 0 x = pi/2 or 3pi/3 so yes. Now you need to find the concavity on (0,pi/2) , (pi/2,3pi/2) , (3pi/2, 2pi)

OpenStudy (campbell_st):

thats correct... so the angles that give than values are in the 3rd and 4th quadrants... so what are the angles...?

OpenStudy (anonymous):

Okay so my points of inflection are pi/2, 3pi/2 and 2pi.

OpenStudy (campbell_st):

2 of them are.... 2 pi isn't a point of inflection

OpenStudy (anonymous):

yes, so what are the interval you need to check the concavity on?

OpenStudy (anonymous):

Oh okay thank you.

OpenStudy (anonymous):

I wish I could give you a medal too sourwing

OpenStudy (campbell_st):

the points of inflection show where the concavity changes... so you well need to look at the intervals left and right of the points of concavity... and the interval between them...

OpenStudy (campbell_st):

the easiest way to do this question is to graph it using some sort of graphing software

OpenStudy (anonymous):

I got -infinity,pi/2 is concave down pi/2, 3pi/2 is concave up and 3pi/2 +infinity is concave down

OpenStudy (campbell_st):

well thats interesting as by looking at the 1st derivative 0 = 1 + 2sin(x) solutions are \[x = \frac{7\pi}{6}...and..... \frac{11\pi}{6}\] and then you use the 2nd derivative test \[f"(\frac{7\pi}{6}) = 2\cos(\frac{7\pi}{6})\] then \[f"(\frac{7\pi}{6}) < 0... a... maximum\] so you have the concavity wrong... you have the curve concave down between the points of inflection.

OpenStudy (campbell_st):

oops you should have the curve concave down between the points of inflection...

OpenStudy (campbell_st):

if you want to use the 2nd derivative... you need to select points either side of the points of inflection and substitute them into the 2nd derivative to test for concavity... hope it helps

OpenStudy (anonymous):

(0,pi/2) concave up (pi/2,3pi/2) concave down (pi/2,2pi) concave up

OpenStudy (campbell_st):

here is the graph of your curve over the domain, finding max and min values from the 1st derivative helps to identify the concavity

OpenStudy (anonymous):

Wow that is very detailed. Thank you for putting the time into it. Okay I see my mistake thank you. I have to go. @campbell_st